# Reassessing quantum-thermodynamic enhancements in continuous thermal machines

### v2

@@ D.3 Three-qubit autonomous refrigerator
for vector $\vec{p}_{G}^{\,\mathrm{cl}}(t)=(\rho_{000},\rho_{001},\rho_{100},\rho_{011},\rho_{111},\rho_{101},\rho_{010})$. In the above equations, we defined $\Gamma_{1}=-(\gamma_{01}^{\mathrm{c}}+\gamma_{01}^{\mathrm{m}}+\gamma_{01}^{\mathrm{h}})$, $\Gamma_{2}=-(\gamma_{01}^{\mathrm{c}}+\gamma_{01}^{\mathrm{m}}+\gamma_{10}^{\mathrm{h}})$, $\Gamma_{3}=-(\gamma_{10}^{\mathrm{c}}+\gamma_{01}^{\mathrm{m}}+\gamma_{01}^{\mathrm{h}})$, $\Gamma_{4}=-(\gamma_{01}^{\mathrm{c}}+\gamma_{10}^{\mathrm{m}}+\gamma_{10}^{\mathrm{h}})$, $\Gamma_{5}=-(\gamma_{10}^{\mathrm{c}}+\gamma_{10}^{\mathrm{m}}+\gamma_{01}^{\mathrm{h}})$, $\Gamma_{6}=-(\gamma_{10}^{\mathrm{c}}+\gamma_{10}^{\mathrm{m}}+\gamma_{10}^{\mathrm{h}})$, $\Gamma_{7}=-(\gamma_{10}^{\mathrm{c}}+\gamma_{01}^{\mathrm{m}}+\gamma_{10}^{\mathrm{h}})$, $\Gamma_{8}=-(\gamma_{01}^{\mathrm{c}}+\gamma_{10}^{\mathrm{m}}+\gamma_{01}^{\mathrm{h}})$ and $\Gamma_{9}=-\frac{1}{2}(\gamma_{d}+\gamma_{2d}+\gamma_{3d}+\gamma_{u}+\gamma_{2u}+\gamma_{3u})$. The corresponding coefficients in the classical analog read:
$\begin{split}&a_{0}^{\mathrm{cl}\,\prime}=-4g^{2}\gamma_{c}\gamma_{h}\gamma_{m}\left(\bar{n}_{h}\left(\bar{n}_{c}+\bar{n}_{m}+1\right)-\bar{n}_{c}\bar{n}_{m}\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}\left(2\bar{n}_{h}+1\right)\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)\\ &\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)\\ &a_{0}^{\mathrm{cl}\,\prime\prime}=-4g^{2}\gamma_{c}\gamma_{h}\gamma_{m}\left(\bar{n}_{h}\left(\bar{n}_{c}\left(2\bar{n}_{m}+1\right)+\bar{n}_{m}+1\right)+\bar{n}_{c}\bar{n}_{m}\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}\left(2\bar{n}_{h}+1\right)\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)\\ &\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)\\ &a_{1}^{\mathrm{cl}\,\prime}=-8g^{2}\gamma_{c}\gamma_{h}\gamma_{m}\left(\bar{n}_{h}\left(\bar{n}_{c}+\bar{n}_{m}+1\right)-\bar{n}_{c}\bar{n}_{m}\right)(3\gamma_{c}\left(2\bar{n}_{c}+1\right)\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)+\gamma_{c}^{2}\left(2\bar{n}_{c}+1\right){}^{2}\\ &+3\gamma_{h}\gamma_{m}\left(2\bar{n}_{h}+1\right)\left(2\bar{n}_{m}+1\right)+\gamma_{h}^{2}\left(2\bar{n}_{h}+1\right){}^{2}+\gamma_{m}^{2}\left(2\bar{n}_{m}+1\right){}^{2}),\\ \end{split}$ (81)
@@
D.4 Noise-induced-coherent machine
Finally, using the Hamiltonian and Lindblad operators from Sec. II.3, we obtain the generic form of $W_{G}$ for the noise-induced-coherence machine. In contrast with the previous cases, now we obtain non-zero real coherence in the steady state between states |$\ket{\alpha}$ and |$\ket{\beta}$(see Appendix C). The relevant elements of the $W_{G}$ matrix then connect the populations of the four levels and the real part of the coherence between states |$\ket{\alpha}$ and |$\ket{\beta}$:
$W_{G}=\left(\begin{array}[]{ccccc}-(\gamma_{0\alpha}+\gamma_{01})&\gamma_{10}~{}\mathrm{e}^{-i\chi_{\mathrm{w}}}&\gamma_{\alpha 0}~{}\mathrm{e}^{-i\chi_{\mathrm{c}}}&0&0\\ \gamma_{01}~{}\mathrm{e}^{i\chi_{\mathrm{w}}}&-(\gamma_{1\alpha}+\gamma_{1\beta}+\gamma_{10})&\gamma_{\alpha 1}~{}\mathrm{e}^{-i\chi_{\mathrm{h}}}&\gamma_{\beta 1}~{}\mathrm{e}^{-i\chi_{\mathrm{h}}}&2\sqrt{\gamma_{\alpha 1}\gamma_{\beta 1}}\\ \gamma_{0\alpha}~{}\mathrm{e}^{i\chi_{\mathrm{c}}}&\gamma_{1\alpha}~{}\mathrm{e}^{i\chi_{\mathrm{h}}}&-(\gamma_{\alpha 1}+\gamma_{\alpha 0})&0&-\sqrt{\gamma_{\alpha 1}\gamma_{\beta 1}}\\ 0&\gamma_{1\beta}~{}\mathrm{e}^{i\chi_{\mathrm{h}}}&0&-\gamma_{\beta 1}&-\sqrt{\gamma_{\alpha 1}\gamma_{\beta 1}}\\ 0&2\sqrt{\gamma_{1\alpha}\gamma_{1\beta}}&-\sqrt{\gamma_{\alpha 1}\gamma_{\beta 1}}&-\sqrt{\gamma_{\alpha 1}\gamma_{\beta 1}}&\gamma_{\alpha 0}+\gamma_{\alpha 1}+\gamma_{\beta 1}\end{array}\right),$
@@ D.4 Noise-induced-coherent machine
$\begin{split}&a_{0}^{\prime}=\gamma_{\text{eg}}\bar{n}_{c}\bar{n}_{h}\left(\bar{n}_{h}-\bar{n}_{c}\right)\left(\gamma_{c}{}^{\alpha}\right){}^{2}\gamma_{h}{}^{\alpha}\gamma_{h}{}^{\beta}\\ &a_{0}^{\prime\prime}=\gamma_{\text{eg}}\bar{n}_{c}\bar{n}_{h}\left(\bar{n}_{c}\left(2\bar{n}_{h}+1\right)+\bar{n}_{h}\right)\left(\gamma_{c}{}^{\alpha}\right){}^{2}\gamma_{h}{}^{\alpha}\gamma_{h}{}^{\beta}\\ &a_{1}=-\gamma_{c}{}^{\alpha}\gamma_{h}{}^{\beta}\left(\gamma_{\text{eg}}\left(\bar{n}_{c}\left(4\bar{n}_{h}+1\right)+\bar{n}_{h}\right)\left(\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\left(\gamma_{h}{}^{\alpha}+\gamma_{h}{}^{\beta}\right)\right)+\bar{n}_{c}\bar{n}_{h}\left(\bar{n}_{c}\left(4\bar{n}_{h}+3\right)+3\bar{n}_{h}+2\right)\gamma_{c}{}^{\alpha}\gamma_{h}{}^{\alpha}\right)\\ &a_{1}^{\prime}=-\gamma_{\text{eg}}\left(\bar{n}_{c}-\bar{n}_{h}\right)\gamma_{c}{}^{\alpha}\gamma_{h}{}^{\alpha}\left(\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\gamma_{h}{}^{\alpha}\right).\\ \end{split}$ (82)
@@
$W_{G}^{\mathrm{cl}}=\left(\begin{array}[]{ccccc}-(\gamma_{0\alpha}+\gamma_{01})&\gamma_{10}~{}\mathrm{e}^{-i\chi_{\mathrm{w}}}&\gamma_{\alpha 0}~{}\mathrm{e}^{-i\chi_{\mathrm{c}}}&0\\ \gamma_{01}~{}\mathrm{e}^{i\chi_{\mathrm{w}}}&-(\gamma_{1\alpha}^{\mathrm{cl}}+\gamma_{1\beta}^{\mathrm{cl}}+\gamma_{10})&\gamma_{\alpha 1}^{\mathrm{cl}}~{}\mathrm{e}^{-i\chi_{\mathrm{h}}}&\gamma_{\beta 1}^{\mathrm{cl}}~{}\mathrm{e}^{-i\chi_{\mathrm{h}}}\\ \gamma_{0\alpha}~{}\mathrm{e}^{i\chi_{\mathrm{c}}}&\gamma_{1\alpha}^{\mathrm{cl}}~{}\mathrm{e}^{i\chi_{\mathrm{h}}}&-(\gamma_{\alpha 1}^{\mathrm{cl}}+\gamma_{\alpha 0}+\gamma_{\alpha\beta}^{\mathrm{cl}})&\gamma_{\alpha\beta}^{\mathrm{cl}}\\ 0&\gamma_{1\beta}^{\mathrm{cl}}~{}\mathrm{e}^{i\chi_{\mathrm{h}}}&\gamma_{\alpha\beta}^{\mathrm{cl}}&-(\gamma_{\beta 1}^{\mathrm{cl}}+\gamma_{\alpha\beta}^{\mathrm{cl}})\end{array}\right),$
@@
$\begin{split}&a_{0}^{\mathrm{cl}\,\prime}=\frac{\gamma_{\text{eg}}\bar{n}_{c}\bar{n}_{h}\left(\bar{n}_{c}-\bar{n}_{h}\right)\left(\gamma_{c}{}^{\alpha}\right){}^{2}\gamma_{h}{}^{\alpha}\gamma_{h}{}^{\beta}}{\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\left(\gamma_{h}{}^{\alpha}+\gamma_{h}{}^{\beta}\right)}\\ &a_{0}^{\mathrm{cl}\,\prime\prime}=\frac{\gamma_{\text{eg}}\bar{n}_{c}\bar{n}_{h}\left(\bar{n}_{c}-\bar{n}_{h}\right)\left(\gamma_{c}{}^{\alpha}\right){}^{2}\gamma_{h}{}^{\alpha}\gamma_{h}{}^{\beta}}{\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\left(\gamma_{h}{}^{\alpha}+\gamma_{h}{}^{\beta}\right)}\\ &a_{1}^{\mathrm{cl}}=\gamma_{c}{}^{\alpha}\gamma_{h}{}^{\beta}\left(\gamma_{\text{eg}}\left(\bar{n}_{c}\left(4\bar{n}_{h}+1\right)+\bar{n}_{h}\right)+\frac{\bar{n}_{c}\bar{n}_{h}\left(\bar{n}_{c}\left(4\bar{n}_{h}+3\right)+3\bar{n}_{h}+2\right)\gamma_{c}{}^{\alpha}\gamma_{h}{}^{\alpha}}{\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\left(\gamma_{h}{}^{\alpha}+\gamma_{h}{}^{\beta}\right)}\right)\\ &a_{1}^{\mathrm{cl}\,\prime}=\frac{\gamma_{\text{eg}}\left(\bar{n}_{c}-\bar{n}_{h}\right)\gamma_{c}{}^{\alpha}\gamma_{h}{}^{\alpha}\left(\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\left(\gamma_{h}{}^{\alpha}-\gamma_{h}{}^{\beta}\right)\right)}{\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\left(\gamma_{h}{}^{\alpha}+\gamma_{h}{}^{\beta}\right)},\\ \end{split}$ (83)
@@
Appendix E Proof of Theorem I
In this appendix we provide the proof of the Theorems 1 enunciated in Sec. V. The proof follows from the analytical evaluation of the sing of the ratio $\mathcal{R}$ in Eq. (31), which can be reduced to the evaluation of the difference between the variances in quanum and classical-equivalent machines, $\mathrm{Var}[\dot{N}^{\mathrm{cl}}]-\mathrm{Var}[\dot{N}]$. Using the explicit expressions for the variances from the Inverse Counting Statistics method (see App. B) provided in Eq. (73), that difference can be written as:
@@ Appendix E Proof of Theorem I
where we took advantage from the fact that, by construction of the classical equivalent, $\langle\dot{N}\rangle=\langle\dot{N}^{\mathrm{cl}}\rangle$. In the following we evaluate the three under-brace terms in Eq. (84) for the inverse counting coefficients obtained for Hamiltonian-induced coherence [Eqs. (75) and (77)]. For the first term $(1)$ we obtain always a vanishing contribution:
@@
which is always positive since the rates are positive and $\mathcal{Z}>0$ is a positive (normalization) factor. Moreover since this term is multiplied by $\langle\dot{N}\rangle^{2}$ in Eq. (84), it always leads to a positive contribution to $\mathrm{Var}[\dot{N}^{\mathrm{cl}}]-\mathrm{Var}[\dot{N}]$.
@@
$\begin{split}&\left(\dfrac{a_{1}^{\prime}}{a_{1}}-\dfrac{a_{1}^{cl\prime}}{a_{1}^{\mathrm{cl}}}\right)=\dfrac{1}{\mathcal{Z}}(\Gamma_{\text{vs}}\left(\Gamma_{\text{ms}}\Gamma_{\text{su}}\gamma_{\text{um}}-\gamma_{\text{mu}}\Gamma_{\text{sm}}\Gamma_{\text{us}}\right)\\ &+\gamma_{\rm{vm}}(\gamma_{\rm{um}}\Gamma_{\rm{ms}}\Gamma_{\rm{su}}-\gamma_{\rm{mu}}\Gamma_{\rm{us}}\Gamma_{\rm{sm}})\\ &+(\gamma_{\rm{mu}}\Gamma_{\rm{us}}\Gamma_{\rm{sv}}\gamma_{\rm{vm}}-\gamma_{\rm{um}}\gamma_{\rm{mv}}\Gamma_{\rm{vs}}\Gamma_{\rm{su}}))\end{split}$ (87)
where differences between the rates associated to certain cycles in opposite directions appear. Notably, however, none of such cycles involve the coherent transition between levels |$\ket{u}$ and |$\ket{v}$. Now we show that the contribution in Eq. (87) is zero in the cases of unicyclic and multicycle (multipartite) quantum thermal machines. First, in the case of unicyclic machines we will (trivially) have $\Gamma_{\rm{us}}=\Gamma_{\rm{su}}=0$ and $\Gamma_{\rm{vm}}=\Gamma_{\rm{mv}}=0$, which makes all the different terms in Eq. (87) equal to zero. Second, we consider a multicycle structure in the thermal machine where all cycles have the same number of contributions from the different baths. This is the case e.g. of multipartite systems under local dissipation. In that case we have that to close cycles that do not include the coherent transition between the levels |$\ket{\rm u}$ and |$\ket{\rm v}$, we will have the same number of quanta in and out from each given bath during the cycle (see e.g. Fig. 2b for an illustration) and therefore there will be no net probability flux in one or the other direction. This property implies that three three parenthesis in Eq. (87) cancel.
@@
which is always greater that zero in out-of-equilibrium scenarios. Since $\mathcal{R}$ is strictly greater than zero just in the case of $\langle\dot{N}\rangle\neq 0$ which corresponds with the out-of-equilibrium scenario.
Appendix F Details on the NIC machine
In this appendix we discuss in more detail the change of variables introduced in Sec. VI.3 to decouple one of the two collective transitions of the NIC coherent machine, together with extra comments regarding the symmetric limit in which all the spontaneous emission rates are equal.
Figure 10: Fluctuations ratio $\mathcal{R}$ (blue solid line) and variance in the quantum system (orange doted line) as functions of the ratio of the original spontaneous emission rates (prior to the change of variables). The system parameters are: $\beta_{\mathrm{w}}\rightarrow 0$, $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{h}}/\beta_{\mathrm{c}}=0.9$, $\epsilon_{2}=5$, $\epsilon_{1}=3$, and $\gamma_{\mathrm{c}}^{(\mathrm{a}/\mathrm{b})}=10^{-3}$ (energies are given in units of $k_{B}T_{\rm c}=1$).

### v4

@@ F.3 Three-qubit autonomous refrigerator
for vector $\vec{p}_{G}^{\,\mathrm{cl}}(t)=(\rho_{000},\rho_{001},\rho_{100},\rho_{011},\rho_{111},\rho_{101},\rho_{010})$. The corresponding coefficients in the classical equivalent machine read:
$a_{0}^{\mathrm{cl}\,\prime}=$ $-4g^{2}\gamma_{c}\gamma_{h}\gamma_{m}\left(\bar{n}_{h}\left(\bar{n}_{c}+\bar{n}_{m}+1\right)-\bar{n}_{c}\bar{n}_{m}\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}\left(2\bar{n}_{h}+1\right)\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)$
$~{}~{}\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right),$
$a_{0}^{\mathrm{cl}\,\prime\prime}=$ $-4g^{2}\gamma_{c}\gamma_{h}\gamma_{m}\left(\bar{n}_{h}\left(\bar{n}_{c}\left(2\bar{n}_{m}+1\right)+\bar{n}_{m}+1\right)+\bar{n}_{c}\bar{n}_{m}\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{h}\left(2\bar{n}_{h}+1\right)\right)\left(\gamma_{c}\left(2\bar{n}_{c}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)$
$~{}\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right),$
$a_{1}^{\mathrm{cl}\,\prime}=$ $-8g^{2}\gamma_{c}\gamma_{h}\gamma_{m}\left(\bar{n}_{h}\left(\bar{n}_{c}+\bar{n}_{m}+1\right)-\bar{n}_{c}\bar{n}_{m}\right)(3\gamma_{c}\left(2\bar{n}_{c}+1\right)\left(\gamma_{h}\left(2\bar{n}_{h}+1\right)+\gamma_{m}\left(2\bar{n}_{m}+1\right)\right)+\gamma_{c}^{2}\left(2\bar{n}_{c}+1\right){}^{2}$
$~{}+3\gamma_{h}\gamma_{m}\left(2\bar{n}_{h}+1\right)\left(2\bar{n}_{m}+1\right)+\gamma_{h}^{2}\left(2\bar{n}_{h}+1\right){}^{2}+\gamma_{m}^{2}\left(2\bar{n}_{m}+1\right)^{2}),$ (82)
@@
F.4 Noise-induced-coherent machine
Finally, using the Hamiltonian and Lindblad operators from Eq. (31) and (32), we obtain the generic form of $W_{G}$ for the noise-induced-coherence machine. In contrast with the previous cases, now we obtain non-zero real coherence in the steady state between states |$\ket{\alpha}$ and |$\ket{\beta}$(see App. D). The relevant elements of the $W_{G}$ matrix then connect the populations of the four levels and the real part of the coherence between states |$\ket{\alpha}$ and |$\ket{\beta}$:
$W_{G}=\left(\begin{array}[]{ccccc}-(\gamma_{0\alpha}+\gamma_{01})&\gamma_{10}~{}\mathrm{e}^{-i\chi}&\gamma_{\alpha 0}&0&0\\ \gamma_{01}~{}\mathrm{e}^{i\chi}&-(\gamma_{1\alpha}+\gamma_{1\beta}+\gamma_{10})&\gamma_{\alpha 1}&\gamma_{\beta 1}&2\sqrt{\gamma_{\alpha 1}\gamma_{\beta 1}}\\ \gamma_{0\alpha}&\gamma_{1\alpha}&-(\gamma_{\alpha 1}+\gamma_{\alpha 0})&0&-\sqrt{\gamma_{\alpha 1}\gamma_{\beta 1}}\\ 0&\gamma_{1\beta}&0&-\gamma_{\beta 1}&-\sqrt{\gamma_{\alpha 1}\gamma_{\beta 1}}\\ 0&2\sqrt{\gamma_{1\alpha}\gamma_{1\beta}}&-\sqrt{\gamma_{\alpha 1}\gamma_{\beta 1}}&-\sqrt{\gamma_{\alpha 1}\gamma_{\beta 1}}&\gamma_{\alpha 0}+\gamma_{\alpha 1}+\gamma_{\beta 1}\end{array}\right),$
@@ F.4 Noise-induced-coherent machine
$\begin{split}&a_{0}^{\prime}=\gamma_{\text{eg}}\bar{n}_{c}\bar{n}_{h}\left(\bar{n}_{h}-\bar{n}_{c}\right)\left(\gamma_{c}{}^{\alpha}\right){}^{2}\gamma_{h}{}^{\alpha}\gamma_{h}{}^{\beta},~{}~{}~{}~{}a_{0}^{\prime\prime}=\gamma_{\text{eg}}\bar{n}_{c}\bar{n}_{h}\left(\bar{n}_{c}\left(2\bar{n}_{h}+1\right)+\bar{n}_{h}\right)\left(\gamma_{c}{}^{\alpha}\right){}^{2}\gamma_{h}{}^{\alpha}\gamma_{h}{}^{\beta},\\ &a_{1}=-\gamma_{c}{}^{\alpha}\gamma_{h}{}^{\beta}\left(\gamma_{\text{eg}}\left(\bar{n}_{c}\left(4\bar{n}_{h}+1\right)+\bar{n}_{h}\right)\left(\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\left(\gamma_{h}{}^{\alpha}+\gamma_{h}{}^{\beta}\right)\right)+\bar{n}_{c}\bar{n}_{h}\left(\bar{n}_{c}\left(4\bar{n}_{h}+3\right)+3\bar{n}_{h}+2\right)\gamma_{c}{}^{\alpha}\gamma_{h}{}^{\alpha}\right),\\ &a_{1}^{\prime}=-\gamma_{\text{eg}}\left(\bar{n}_{c}-\bar{n}_{h}\right)\gamma_{c}{}^{\alpha}\gamma_{h}{}^{\alpha}\left(\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\gamma_{h}{}^{\alpha}\right).\end{split}$ (83)
@@
$W_{G}^{\mathrm{cl}}=\left(\begin{array}[]{ccccc}-(\gamma_{0\alpha}+\gamma_{01})&\gamma_{10}~{}\mathrm{e}^{-i\chi}&\gamma_{\alpha 0}&0\\ \gamma_{01}~{}\mathrm{e}^{i\chi}&-(\gamma_{1\alpha}^{\mathrm{cl}}+\gamma_{1\beta}^{\mathrm{cl}}+\gamma_{10})&\gamma_{\alpha 1}^{\mathrm{cl}}&\gamma_{\beta 1}^{\mathrm{cl}}\\ \gamma_{0\alpha}&\gamma_{1\alpha}^{\mathrm{cl}}&-(\gamma_{\alpha 1}^{\mathrm{cl}}+\gamma_{\alpha 0}+\gamma_{\alpha\beta}^{\mathrm{cl}})&\gamma_{\alpha\beta}^{\mathrm{cl}}\\ 0&\gamma_{1\beta}^{\mathrm{cl}}&\gamma_{\alpha\beta}^{\mathrm{cl}}&-(\gamma_{\beta 1}^{\mathrm{cl}}+\gamma_{\alpha\beta}^{\mathrm{cl}})\end{array}\right),$
@@
$\begin{split}&a_{0}^{\mathrm{cl}\,\prime}=\frac{\gamma_{\text{eg}}\bar{n}_{c}\bar{n}_{h}\left(\bar{n}_{c}-\bar{n}_{h}\right)\left(\gamma_{c}{}^{\alpha}\right){}^{2}\gamma_{h}{}^{\alpha}\gamma_{h}{}^{\beta}}{\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\left(\gamma_{h}{}^{\alpha}+\gamma_{h}{}^{\beta}\right)},~{}~{}~{}~{}a_{0}^{\mathrm{cl}\,\prime\prime}=\frac{\gamma_{\text{eg}}\bar{n}_{c}\bar{n}_{h}\left(\bar{n}_{c}-\bar{n}_{h}\right)\left(\gamma_{c}{}^{\alpha}\right){}^{2}\gamma_{h}{}^{\alpha}\gamma_{h}{}^{\beta}}{\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\left(\gamma_{h}{}^{\alpha}+\gamma_{h}{}^{\beta}\right)},\\ &a_{1}^{\mathrm{cl}}=\gamma_{c}{}^{\alpha}\gamma_{h}{}^{\beta}\left(\gamma_{\text{eg}}\left(\bar{n}_{c}\left(4\bar{n}_{h}+1\right)+\bar{n}_{h}\right)+\frac{\bar{n}_{c}\bar{n}_{h}\left(\bar{n}_{c}\left(4\bar{n}_{h}+3\right)+3\bar{n}_{h}+2\right)\gamma_{c}{}^{\alpha}\gamma_{h}{}^{\alpha}}{\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\left(\gamma_{h}{}^{\alpha}+\gamma_{h}{}^{\beta}\right)}\right),\\ &a_{1}^{\mathrm{cl}\,\prime}=\frac{\gamma_{\text{eg}}\left(\bar{n}_{c}-\bar{n}_{h}\right)\gamma_{c}{}^{\alpha}\gamma_{h}{}^{\alpha}\left(\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\left(\gamma_{h}{}^{\alpha}-\gamma_{h}{}^{\beta}\right)\right)}{\bar{n}_{c}\gamma_{c}{}^{\alpha}+\bar{n}_{h}\left(\gamma_{h}{}^{\alpha}+\gamma_{h}{}^{\beta}\right)},\end{split}$ (84)
@@
Appendix G Proof of Theorem I
In this Appendix we provide the proof of Theorem 1 enunciated in the main text. The proof follows from the analytical evaluation of the sing of the ratio $\mathcal{R}$ in Eq. (15), which can be reduced to the evaluation of the difference between the variances in quantum and classical-equivalent machines, $\mathrm{Var}[\dot{N}^{\mathrm{cl}}]-\mathrm{Var}[\dot{N}]$. Using the explicit expressions for the variances from the Inverse Counting Statistics method (see App. E) provided in Eq. (74), that difference can be written as:
@@ Appendix G Proof of Theorem I
where we took advantage from the fact that, by construction of the classical equivalent, $\langle\dot{N}\rangle=\langle\dot{N}^{\mathrm{cl}}\rangle$. In the following, we evaluate the three under-brace terms in Eq. (85) for the inverse counting coefficients obtained for Hamiltonian-induced coherence [Eqs. (F.1) and (78)]. We first consider the case of resonant driving $(\Delta_{\rm d}=0)$ and then extend the analysis to non-zero detuning. For the first term $(1)$ we obtain always a vanishing contribution:
@@
which is always positive since the rates are positive and $\mathcal{Z}>0$ is a positive (normalization) factor. Moreover, since this term is multiplied by $\langle\dot{N}\rangle^{2}$ in Eq. (85), it always leads to a positive contribution to $\mathrm{Var}[\dot{N}^{\mathrm{cl}}]-\mathrm{Var}[\dot{N}]$.
@@
$\begin{split}\left(\dfrac{a_{1}^{\prime}}{a_{1}}-\dfrac{a_{1}^{cl\prime}}{a_{1}^{\mathrm{cl}}}\right)&=\dfrac{1}{\mathcal{Z}}[\Gamma_{\text{vs}}\left(\Gamma_{\text{ms}}\Gamma_{\text{su}}\gamma_{\text{um}}-\gamma_{\text{mu}}\Gamma_{\text{us}}\Gamma_{\text{sm}}\right)\\ &+\gamma_{\rm{vm}}(\gamma_{\rm{um}}\Gamma_{\rm{ms}}\Gamma_{\rm{su}}-\Gamma_{\rm{us}}\Gamma_{\rm{sm}}\gamma_{\rm{mu}})\\ &+(\gamma_{\rm{mu}}\Gamma_{\rm{us}}\Gamma_{\rm{sv}}\gamma_{\rm{vm}}-\gamma_{\rm{mv}}\Gamma_{\rm{vs}}\Gamma_{\rm{su}}\gamma_{\rm{um}})]\end{split}$ (88)
where differences between the rates associated to certain cycles in opposite directions appear. However, in particular, none of these cycles involves the coherent transition between levels |$\ket{u}$ and |$\ket{v}$. Therefore, for unicyclic machines, we (trivially) have $\Gamma_{\rm{us}}=\Gamma_{\rm{su}}=0$ and $\Gamma_{\rm{vm}}=\Gamma_{\rm{mv}}=0$, which makes all the different terms in Eq. (88) equal to zero. Moreover, this term is also zero for symmetric multicycle quantum thermal machines, where all cycles have the same number of contributions from the different baths. This is the case, e.g. of multipartite systems under local dissipation. In that case we have that to close cycles that do not include the coherent transition between the levels |$\ket{\rm u}$ and |$\ket{\rm v}$, we will have the same number of quanta in and out from each given bath during the cycle (see e.g. Fig. 2b for an illustration) and therefore there will be no net probability flux in one or the other direction. This property implies that each of the three parenthesis in Eq. (88), which are proportional to the conditional probabilities to close three of such cycles, cancels out.
@@
which is always strictly greater that zero in out-of-equilibrium scenarios. The reason is that $\mathcal{R}$ becomes zero only in the case of $\langle\dot{N}\rangle=0$, which corresponds to the equilibrium point.
To analyze the effect of detuning in the driving, we now repeat the same steps as in the earlier demonstration but using the more general expressions derived within this scenario. We find that, as before, (86) does not contribute to $\mathcal{R}$. However, in this case, (87) and (88) do not consistently exhibit a well-defined positive sign in all cases. By considering the types of uni-cyclic and multi-cyclic machines to which our theorem applies, it can be shown that the fluctuation ratio takes the form
$\mathcal{R}=\left(\left[\dfrac{1}{2}\sum_{i}(\gamma_{\rm ui}+\gamma_{\rm vi})\right]^{2}-\Delta_{\rm d}^{2}\right)\Lambda,$ (90)
where $\Lambda\geq 0$ represents a sum of rates (for its explicit expression see Ref. [92]). As a consequence, we have $\mathcal{R}\geq 0$ whenever the expression within parentheses in Eq. (90). Theorem 1 remains thus valid in the presence of detuning when $|\Delta_{\rm d}|<\dfrac{1}{2}\sum_{i}(\gamma_{\rm ui}+\gamma_{\rm vi})$ is verified.
Appendix H Details on the NIC machine analysis
In this Appendix we discuss in more detail the change of variables introduced in Sec. V.3 to decouple one of the two collective transitions of the NIC coherent machine, together with extra comments regarding the symmetric limit in which all the spontaneous emission rates are equal.
Figure 7: Fluctuations ratio $\mathcal{R}$ (blue solid line) for the NIC machine and variance in the quantum system (orange doted line) as functions of the ratio of the original spontaneous emission rates (prior to the change of variables). The system parameters are: $\beta_{\mathrm{w}}\rightarrow 0$, $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{h}}/\beta_{\mathrm{c}}=0.9$, $\epsilon_{2}=5$, $\epsilon_{1}=3$, and $\gamma_{\mathrm{c}}^{(\mathrm{a}/\mathrm{b})}=10^{-3}$ (energies are given in units of $k_{B}T_{\rm c}=1$).